### v1

# Joint Estimation of Identity Verification and Relative Pose for Partial Fingerprints

Abstract
Currently, portable electronic devices are becoming more and more popular. For lightweight considerations, their fingerprint recognition modules usually use limited-size sensors. However, partial fingerprints have few matchable features, especially when there are differences in finger pressing posture or image quality, which makes partial fingerprint verification challenging. Most existing methods regard fingerprint position rectification and identity verification as independent tasks, ignoring the coupling relationship between them — relative pose estimation typically relies on paired features as anchors, and authentication accuracy tends to improve with more precise pose alignment. Consequently, in this paper we propose a method that jointly estimates identity verification and relative pose for partial fingerprints, aiming to leverage their inherent correlation to improve each other. To achieve this, we propose a multi-task CNN (Convolutional Neural Network)-Transformer hybrid network, and design a pre-training task to enhance the feature extraction capability. Experiments on multiple public datasets (NIST SD14, FVC2002 DB1_A & DB3_A, FVC2004 DB1_A & DB2_A, FVC2006 DB1_A) and an in-house dataset show that our method achieves state-of-the-art performance in both partial fingerprint verification and relative pose estimation, while being more efficient than previous methods.
Index Terms:
Fingerprint recognition, partial fingerprint, fingerprint verification, fingerprint pose estimation, transformer.
I Introduction
Fingerprints are unique patterns composed of ridges and valleys on the finger surface. Due to its easy collection, high stability and strong recognizability, this biological characteristic exhibits significant application value. As early as the nineteenth century, researchers had conducted systematic studies on fingerprints [1, 2, 3]. The reliability of fingerprint recognition gained official recognition in the twentieth century and found extensive application within the judicial sphere [3]. With the advancement of devices and algorithms, fingerprint recognition technology has been swiftly applied in civilian and commercial fields, such as electronic devices, financial services, and access control systems, providing people with convenience and security [4].
Fueled by consumer demand and advances in technology, the trend towards integration, miniaturization, and portability has become prominent within the consumer electronics sector. This progression compels manufacturers to consistently reduce the dimensions of sensors to accommodate progressively compact device architectures. Fingerprint sensors, currently one of the most popular biometric modules, have also introduced various miniaturized solutions for extensive use [4]. However, the size of fingerprint sensors is inevitably limited, resulting in a significant reduction in amount of information available for matching [5, 6, 7]. Fig. 1 shows three representative challenging scenarios in partial fingerprint recognition, which can be even marked on fingers with large posture differences or poor quality (such as wear and wrinkles), leading to a substantial degradation in the performance of feature extraction and matching [8]. This not only reduces the user experience, but also poses great security risks [9].
Figure 1: Partial fingerprint recognition faces three challenges: sparse minutiae distribution, similar local texture and significant modal difference. This figure shows corresponding representative examples: (a) Three weakly-textured fingerprints that lack appropriate landmark feature points. (b) Local ridge patterns of fingerprints from different fingers have high similarity. (c) Fingerprint features in different modalities (in this case from left to right are dry, normal and wet respectively) may be missed (green) or incorrectly extracted (red). Visualized feature points are minutiae extracted by VeriFinger [10].
A natural idea for partial fingerprint matching is to transfer existing solutions for rolled or plain fingerprint recognition. Nevertheless, conventional algorithms usually rely on a sufficient number of key points to establish and compare spatial structural relationships [11, 12]. While certain studies use neural networks to extract region descriptors [13, 14, 15], consequently reducing the dependence on point features, these approaches require either a considerable effective area or exceedingly precise alignment. A series of fusion strategies are subsequently proposed at feature [16, 17, 18], image [19, 20] and decision levels [21, 22, 23], but most of them still depend on single-to-single matching or require several densely distributed high-quality images. In recent years, this field has garnered significant interest among researchers, leading to the development of several verification methods [24, 25, 8]. However, to the best of our knowledge, almost all existing fingerprint recoginition methods (compared in Fig. 2) regard pose rectification and identity verification as independent tasks, ignoring the coupling relationship between them. Specifically, the estimation of relative pose typically relies on paired features as anchors, and the accuracy of authentication generally increases with enhanced precision in pose alignment. This is similar to a pictorial jigsaw puzzle where joint decision is suggested about whether (are they neighboring image fragments? ) and how (what’s the reassemble solution? ) any two pieces fit together [26, 27, 28], in order to better balance robustness and speed in the overall process, rather than being made individually.
Figure 2: Frameworks of different fingerprint recognition algorithms. The process passed by parallel arrows indicates that the corresponding modules share weights and their functions can be executed in parallel. The ‘Information Interaction’ extracts and compares features from paired data rather than executing independently and separately. It should be noted that the pose predictor in (a)(c)(d) and (b) estimate the relative and absolute pose respectively.
In this paper, we proposed a multi-task CNN-Transformer hybrid network which jointly estimates identity verification and relative pose for partial fingerprints, called Partial Fingerprint Jigsaw Puzzle Network (FingerPuzzle). The intrinsic connection between feature correspondence and position correlation is exploited, aiming to promote each other to achieve higher performance. Naturally, the efficiency is also improved because these two tasks tasks that are independent in previous methods are compactly integrated. Inspired by works on natural images [29, 30, 31, 32, 33], we utilize a shared weight convolutional network to extract features of local patches and transformers to capture both local and global correlations. Besides, a fingerprint enhancement pre-training task is specially designed to enhance the compatibility with different image modalities and various texture patterns.
Extensive experiments are conducted on abundant datasets, including several public datasets (NIST SD14 [34], FVC2002 DB1_A & DB3_A [35],FVC2004 DB1_A & DB2_A [36], FVC2006 DB1_A [37]) and an in-house dataset. Experimental results show that the proposed algorithm outperforms previous state-of-the-art methods in both partial fingerprint verification and pose estimation, and also demonstrates high efficiency.
The main contributions of our paper can be summarized as:
  • We propose a novel approach for partial fingerprint recognition by jointly estimating the identity verification and relative pose instead of the previous independent stages.
  • A CNN-Transformer hybrid network structure is presented to aggregate their respective advantages in local feature extraction and global information interaction.
  • A lightweight pre-training task on fingerprint enhancement is specifically designed to further improve the generalization ability of our proposed network.
  • Sufficient experiments are conducted on diverse datasets to comprehensively evaluate representative algorithms in partial fingerprints at different sizes, which demonstrates the superiority of our proposed method.
The paper is organized as follows: Section II reviews the related works. Section III introduces the proposed partial fingerprint recognition algorithm. Section IV describes the details and usage of datasets. Section V presents the experimental results and discussions. Finally, we draw conclusions in Section VI.
II Related Work
According to different execution frameworks, existing single-to-single fingerprint recognition algorithms can be mainly divided into three categories: recognition based on key points, recognition based on region descriptors and verification based on pairwise correlation. We call the first two ‘recognition’ because highly generalized features are stored and compared here (in other words, each fingerprint only needs to extract features once) to enable efficient searches in large libraries. On the other hand, ‘verification’ uses richer information (features need to be re-extracted for each match) and is more suitable for confirmation scenarios with high precision requirements, such as unlocking a mobile phone. Fig. 2 shows the corresponding schematic diagrams of these processes.
II-A Recognition Based on Key Points
Minutiae are the most popular features used for fingerprint matching [4]. Researchers define the features based on the attributes of minutiae themselves [11, 12] and auxiliary information (such as orientation [38], period map [39], ridge [40] etc.). Considering the lack of high-quality manually defined minutiae in latent fingerprints, Cao et al.[41] uniformly sampled images into small patches and used a network to extract fixed length vector as virtual minutiae. However, as explained in Fig. 1, these features are not discriminative enough in partial fingerprints n the other hand, some studies introduced universal key points in computer vision field into fingerprints [42, 7]. Features extracted in this way are more densely distributed, but at the same time more sensitive to details and therefore easily confused. There are also algorithms that exploit Level-3 features from high-resolution images [43, 44, 45], which are relatively expensive to deploy for civilian equipment. After getting satisfactory point sets, correspondence between them will be established based on the similarity of features and spatial structures, and then used to calculate relative poses and matching scores.
II-B Recognition Based on Region Representation
Early studies use hand-crafted statistics to characterize fingerprint regions and convert them into fixed length vectors [46, 47, 48, 49]. This feature form significantly improves the speed of indexing at the expense of certain accuracy, and is usually used for rough classification. Engelsma et al.[13] used an end-to-end network to extract fixed-length global representations of minutiae and textures, which greatly improved the performance. Inspired by this, researchers have conducted more exploration and improvement of fixed-length representation methods based on deep learning [14, 50, 51, 15]. These algorithms usually require pose rectification in advance, such as adaptively adjusting the image geometry through a spatial transformation layer [13, 15], using a separate pose estimation network to rectify the absolute finger pose [50, 51], or exploiting the positioning relationships implicit in minutiae [14]. The error of pose estimation is tolerated to a certain extent because the extracted features are a generalization of regional characteristics. However, current methods struggle to address scenarios with incomplete or significantly different regions.
II-C Verification Based on Pairwise Correlation
Compared with key points or fixed-length descriptors, gray-level information of pixels contain richer and more discriminatory information about local textures [52]. Traditional algorithms use image correlation for fine-grained verification, which exhibits a certain reliability [52, 53, 54]. Some works on fingerprint distortion further increase the tolerance of ridge misalignment [55, 56]. With the development of deep learning, researchers apply neural networks to directly extract and verify features between each other, significantly improving the performance and efficiency [57, 24, 25, 58]. He et al.[8] use multi-rotation and multi-size cropped image pairs as multi-inputs to help the network better understand scenarios with large spatial transformations or small overlapping areas. It should be noted that a significant difference between correlation based verification and region descriptor based recognition is that the relative pose of a certain fingerprint is rectified instead of the respective absolute pose. This is because features of such methods are compared more precisely with spatial correspondence thus have stricter requirements for alignment. Obviously, the verification performance will significantly declines with inaccurate pose estimation [8].
Figure 3: An overview of FingerPuzzle. Bars are presented on the left to indicate each phase, where the color definition refers to Fig. 2. The process passed by parallel arrows indicates that the corresponding modules share weights and their functions can be executed in parallel. Dotted parts are only used to assist training and could be pruned in practical verification tasks. The outputs ‘p’ and ‘H’ of respective task heads correspond to the classification probability (whether the input fingerprints come from the same finger) and rigid transformation parameters (relative translation and rotation) respectively.
III Method
In this paper, we propose a CNN-Transformer hybrid network that jointly estimates identity verification and relative pose of partial fingerprint pairs. Fig. 3 gives the complete flowchart of the proposed algorithm. The overall process can be divided into four phases, which are indicated by different color bars in schematic: feature extraction (red), information interaction (gold), identity verification (blue) and pose prediction (green). For an input fingerprint pair, our algorithm first extracts the respective features in parallel using convolutional blocks with shared weights (in stem and stages 1, 2). These two sets of features are then concatenated and fed to transformer blocks in stage 3 (along with some extra microprocessing) for sufficient information interaction both locally and globally. Finally, respective output layers utilize the aggregated information to predict classification probabilities for authentication (stage 4 - C and classifier) and regression values for relative poses (stage 4 - R and regressor). It should be noted that these two tasks are jointly estimated in our proposed network instead of being performed independently as previous methods because we hope to utilize the mutually beneficial coupling relationship between each other, inspired by works on jigsaw puzzles [26, 27, 28]. In particular, pose-related regression layers could be pruned to improve efficiency since only identity identification is necessary in most practices.
III-A Feature Extraction
Some studies have proven that early stacked convolution blocks can significantly improve the stability and performance of subsequent vision transformers [59, 60]. Moreover, the hybrid strategy of using early CNN and subsequent transformer to aggregate local details and global information has achieved great success in multiple fields [29, 30, 31, 32, 33]. Inspired by these, we also use convolution blocks to parallelly extract independent features from each input partial fingerprint in the early stages of proposed network.
Specifically, the feature extraction module sequentially contains: (1) a $3\times 3$ convolutional stem with stride $1$ (output $32$ channels); (2) stacked encoder blocks in stage 2 for preliminary extraction of robust features, which consisting of two Feature Extraction (FE) blocks and a convolutional layer with stride $2$ for downsampling while doubling the channels ( distinguished other convolutions by *); (3) stacked residual blocks in stage 3 to further enhance the expressive ability of neural network while mitigating gradient problems in deep layers. In order to improve robustness in different image qualities, we introduce FE block from a simple but efficient baseline for image restoration tasks [61], which is formally similar to the combination of Mobile Convolution (MBConv) block [62] and Multilayer Perceptron (MLP) block [63] with some minor adjustments. The details of corresponding architecture is presented on the left side of Fig. 3, where ‘DConv’ and ‘CA’ are depthwise convolution [62] and channel attention [64] respectively.
Furthermore, we designed a pre-training task and corresponding data augmentation strategies to help the network better understand and refine the essential features of fingerprints. As shown in Fig. 4, the pre-training network consists only of FE block and convolution without additional complex components, arranged in a U-shape for fingerprint enhancement follows [23, 65]. For data preparation of inputs, we randomly crop small patches ($128\times 128$ in this paper) from high-quality rolled fingerprints as prototypes to ensure sufficient texture patterns at each position while accelerating the training process. The representative examples of corresponding data augmentation are given on the right of Fig. 4, including:
  • Mirror flip, rotation (by 90, 180 or 270 degree) and grayscale inversion to increase data diversity;
  • Gaussian blur and noise with different cores and variances to simulate irregular sensor noise;
  • Dilation and erosion with structural elements of different sizes to simulate dry and wet fingers respectively;
  • While blobs with random parameters which is added or subtracted to simulate fingerprint-specific noise in incomplete or contaminated scenarios [66].
These augmentation strategies are randomly selected and combined to generate the final input image. On the other hand, the corresponding binary fingerprint calculated by VeriFinger SDK 12.0 [10] is used as the target of this enhancement task. The parameters of stem and stage 1 are then loaded into corresponding modules of the main network.
Fig. 5 shows representative examples of several typical enhancement methods. It can be seen that our approach is smoother than methods based on grayscale mapping [67] or filtering [68]. In addition, our approach also shows certain robust performance when regional information is missing or wrong. This qualitative comparison strongly proves that our pre-training strategy can help the network understand more essential and complete features, and also avoids additional cumbersome enhancement steps. Further quantitative results are given in ablation experiments below.
Figure 4: Illustration of fingerprint enhancement task for pre-training. The network architecture is shown on the left, where solid and dotted boxes indicate the pre-trained parameters will (or not) be loaded into corresponding modules in Fig. 3. The right subfigure gives representative examples of inputs (original & augmented image) and targets (enhanced image) generation.
Figure 5: Comparison of representation enhancement methods on four low-quality fingerprints. Each column from left to right is original image and corresponding enhancement results of CLAHE [67], FingerNet [68], VeriFinger [10] and our proposed method.
III-B Information Interaction
Inspired by works on natural images [29, 30, 31, 32, 33], we utilize transformers to capture long-range dependencies within and across paired features, thereby ensuring sufficient information interaction. The paired features $f_{1}$ and $f_{2}$ are first biased with a sinusoidal positional encoding term [29], and then spatially flattened into 1D sequences and fed into stacked transformers with interleaved self- and cross- attention following [69, 33]. Let $x$ and $y$ denote the 2D position, $i$ and $d$ denote the current and total dimensions, the additional absolute position information is calculated as:
$PE_{(x,y,4i)}=\sin\left(x/10000^{4i/d}\right)\;,$ (1)
$PE_{(x,y,4i+1)}=\cos\left(x/10000^{4i/d}\right)\;,$
$PE_{(x,y,4i+2)}=\sin\left(y/10000^{4i/d}\right)\;,$
$PE_{(x,y,4i+3)}=\cos\left(y/10000^{4i/d}\right)\;.$
It should be noted that following two attention blocks have the same structure, including residual-connected Multi-Head Attention (MHA) and Feed-Forward Network (FFN) [70], only the data streams are differentiated by their respective expected effects. Specifically, let $\operatorname{SA}\left(f_{i},f_{j}\right)$ (Self-Attention) and $\operatorname{CA}\left(f_{i},f_{j}\right)$ (Cross-Attention) represent the corresponding sub-blocks in Fig. 3, the complete process of transformer block is defined as:
$f_{1}^{\mathrm{S}}=\operatorname{SA}\left(f_{1},f_{1}\right)$ $\;,\;f_{2}^{\mathrm{S}}=\operatorname{SA}\left(f_{2},f_{2}\right)\;,$ (2)
$f_{1}^{\mathrm{C}}=\operatorname{CA}\left(f_{1}^{\mathrm{S}},f_{2}^{\mathrm{S}}\right)$ $\;,\;f_{2}^{\mathrm{C}}=\operatorname{CA}\left(f_{2}^{\mathrm{S}},f_{1}^{\mathrm{S}}\right)\;.$
The linear projection of $f_{i}$ and $f_{j}$ in MHA from left to right are $Query$, $Key$ and $Value$ matrices, which are then used for dot-product attention calculation formally as:
$\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V\;,$ (3)
where $d_{k}$ is the $Key$ dimention and the number of heads is fixed to $6$ at this stage. In order to further refine features while reducing information loss, a patch merging layer [71] is subsequently incorporated for downsampling. Finally, the positional encoding is applied again in the same way of Equation 1.
III-C Identification Verification & Pose Prediction
Similar to the previous phase, features are separately fed into stacked identical transformer blocks with head of $12$ to prepare for the corresponding task head. We define identity verification as a binary classification task, that is, judging whether the input image pairs come from a same finger. Let $\mathcal{F}$, $\operatorname{BN}$, $\operatorname{Swish}$ and $\operatorname{AvgPool}$ denote fully connected layer, batch normalization, swish activation and adaptive average pooling respectively, the classification probability $p$ is calculated as:
$\operatorname{MLP}[\cdot]$ =$=\left\{\mathcal{F},\operatorname{BN},\operatorname{Swish}\;,\mathcal{F},\operatorname{BN},\operatorname{Swish}\right\}\;,$ (4)
$\operatorname{Cla}[\cdot]$ =$=\left\{\mathcal{F},\operatorname{AvgPool},\operatorname{MLP}\right\}\;,$
$p$ $=\operatorname{Sigmoid}(\operatorname{Cla}[f])\;,$
[…]
$\operatorname{Reg}[\cdot]$ =$=\left\{\mathcal{F},\operatorname{AvgPool},\operatorname{MLP}\right\}\;,$ (5)
$H$ $=\operatorname{Reg}[f]\;.$
Following previous work [8], we represent the rotation relationship based on relative values of sine and cosine to circumvent the constraints on numerical range of angles, and directly describe the translation in terms of horizontal and vertical displacements. With the prediction target of $H=[r_{\mathrm{c}},r_{\mathrm{s}},t_{\mathrm{x}},t_{\mathrm{y}}]$, the rigid alignment relationship of paired fingerprints can be modeled as:
$\cos\theta={r_{\mathrm{c}}}/{\sqrt{r_{\mathrm{c}}^{2}+r_{\mathrm{s}}^{2}}}\;,\;\sin\theta={r_{\mathrm{s}}}/{\sqrt{r_{\mathrm{c}}^{2}+r_{\mathrm{s}}^{2}}}\;,$ (6)
$\left[\begin{array}[]{l}u_{\mathrm{a}}\\ v_{\mathrm{a}}\\ 1\end{array}\right]=\left[\begin{array}[]{ccc}\cos\theta&-\sin\theta&t_{\mathrm{x}}\\ \sin\theta&\cos\theta&t_{\mathrm{y}}\\ 0&0&1\end{array}\right]\cdot\left[\begin{array}[]{l}u_{\mathrm{b}}\\ v_{\mathrm{b}}\\ 1\end{array}\right]\;,$
where $\theta$ is the relative rotation angle, $(u,v)$ represents any corresponding position in images $1$ and $2$.
III-D Loss Function
We optimize our network parameters through a comprehensive objective function consisting of two items. The total loss can be formulated as follows:
$\mathcal{L}=\mathcal{L}_{\mathrm{cla}}+\lambda\cdot\mathcal{L}_{\mathrm{reg}}\;,$ (7)
where $\mathcal{L}_{\mathrm{cla}}$ and $\mathcal{L}_{\mathrm{reg}}$ are the identity classification term and pose regression term, respectively. The trade-off parameter $\lambda$ is set to $0.002$ in this paper. Additionally, an extra image segmentation loss $\mathcal{L}_{\mathrm{seg}}$ is proposed for fingerprint enhancement, which is applied only during pre-training.
III-D1 Identity Verification Loss
For classification of identity verification, we adapt focal loss [72] to help our model focus on difficult samples:
$\mathcal{L}_{\mathrm{cla}}=\left\{\begin{array}[]{c}-\alpha\left(1-p\right)^{\gamma}\log p,\text{ if }y=1\\ -\left(1-\alpha\right)p^{\gamma}\log(1-p),\text{ if }y=0\end{array}\right.\;,$ (8)
where $p$ represents the predicted probability, $y$ represents the 0-1 label of ground truth. Balance factors $\alpha$ and $\gamma$ are set to $0.2$ and $2.0$ empirically.
III-D2 Pose Prediction Loss
For regression of pose prediction, weighted mean square error is utilzed for supervision:
$\mathcal{L}_{\mathrm{reg}}=$ $\omega\cdot\left[(\cos\theta^{*}-\cos\theta)^{2}+(\sin\theta^{*}-\sin\theta)^{2}\right]$ (9)
$+\left(1-\omega\right)\cdot\left[\left(t^{*}_{\mathrm{x}}-t_{\mathrm{x}}\right)^{2}+\left(t^{*}_{\mathrm{y}}-t_{\mathrm{y}}\right)^{2}\right]$
where the superscript * indicates the ground truth, otherwise the estimated result. The relative angle $\theta$ and pixel level displacement $t$ is converted from the originally defined output form $H$ and Equation 6. Same as [8], we fix $\omega$ to $0.99$ to balance the optimization of rotation and translation.
III-D3 Image Segmentation Loss
We enhance the input fingerprints in form of binary segmentation, where the grayscale differences on modalities are minimized to make the texture pattern of ridges more prominent. A simplified focal loss [72] is defined as the corresponding metric:
$\mathcal{L}_{\mathrm{seg}}$ $=-\frac{1}{|M|}\sum_{M}\left(1-q\right)^{\gamma}\log\left(q\right),$ (10)
$q$ $=y\cdot p+\left(1-y\right)\cdot\left(1-p\right),$
where $p$ is the probability of foreground, $y$ is the ground truth of corresponding binary image obtained by VeriFinger [10], and $M$ is the image mask. The hyperparameter $\gamma$ is fixed to $2.0$ in experiments.
IV Dataset Description
Comprehensive evaluations are conducted on extensive public datasets, including the representative benchmarks NIST SD14 [34], FVC2002 DB1_A & DB3_A [35], FVC2004 DB1_A & DB2_A [36], and FVC2006 DB2_A [37]. In addition, we also use a private partial fingerprint dataset THU Small to further examine the performance in smartphone situations. Table I presents a detailed description about the composition and usage of these datasets. In particular, we merge several fingerprint databases into a larger dataset, called the Hybrid Dataset (Hybrid DB) , to make the implementation of various experiments more convenient and sufficient. On the other hand, we further evaluate the performance of cross-domain generalization on other separate datasets. Rolled fingerprint is utilized because it covers a more complete area at different pressing poses than plain fingerprints. Image examples of different datasets are shown in Fig. 6. It can be seen that datasets selected in our experiments contains diverse fingerprint impressions with different modalities (normal, dry, wet, polluted, wrinkled, etc.).
TABLE I: All fingerprint datasets used in experiments
Type Dataset Scanners Technolohy Image Description Usage
Rolled NIST SD14 a[34] Inking Fingerprint Card $832\times 768$ 27,000 fingers $\times$ 2 impressions train & test c,h
Plain THU Small b Capacitive Smartphone $160\times 160$ 100 fingers $\times$ 8 impressions test d
FVC2002 DB1_A [35] Optical Identix $388\times 374$ 100 fingers $\times$ 8 impressions test e
FVC2002 DB3_A [35] Capacitive Precise Biometrics $300\times 300$ 100 fingers $\times$ 8 impressions test e
FVC2004 DB1_A [36] Optical CrossMatch $640\times 480$ 100 fingers $\times$ 8 impressions train & test f,h
FVC2004 DB2_A [36] Optical Digital Persona $328\times 364$ 100 fingers $\times$ 8 impressions train & test f,h
FVC2006 DB2_A [37] Optical BiometriKa $400\times 560$ 140 fingers $\times$ 12 impressions train & test g,h
  • a
    NIST SD14 was publicly available but later removed from public domain by NIST.
  • b
    Corresponding dataset is private.
  • c
    A total of $108,000$ genuine match pairs and $108,000$ imposter match pairs are used in experiments.
  • d
    A total of $5,600$ genuine match pairs and $9,900$ imposter match pairs are used in experiments.
  • e
    A total of $5,600$ genuine match pairs and $5,600$ imposter match pairs are used in experiments.
  • f
    A total of $22,400$ genuine match pairs and $22,400$ imposter match pairs are used in experiments.
  • g
    A total of $73,920$ genuine match pairs and $73,920$ imposter match pairs are used in experiments.
  • h
    Datasets are merged and proportionally divided for training ($95\,\%$, called Hybrid DB_A) and testing ($5\,\%$, called Hybrid DB_B). Identities between these two subsets are completely isolated.
Figure 6: Image examples from different fingerprint datasets (a) NIST SD14 [34], (b) THU Small, (c) FVC2002 DB1_A [35], (d) FVC2002 DB3_A [35], (e) FVC2004 DB1_A [36], (f) FVC2004 DB2_A [36], (g) FVC2006 DB2_A [37].
IV-A Parital Fingerprint Pair Simulation
Considering the lack of publicly available partial fingerprints, image patches of specified sizes are synthesized from corresponding rolled or plain datasets, following previous methods [8, 22]. For any two fingerprints from the same finger, we first use VeriFinger [10] to align them based on matching minutiae. Alignment in this way is accurate in most cases because selected original image pairs generally have large enough overlap areas of high quality. Subsequently, we calculated the common mask of aligned fingerprints and randomly select the patch center of one partial fingerprint from it. Taking this point as the center, another patch center is uniformly sampled in polar coordinates with a random angle ranging from $-180^{\circ}$ to $180^{\circ}$ and random radius ranging from 0 pixel to 100 pixels. Finally, full fingerprints are cropped into partial images of $160\times 160$, $128\times 128$ and $96\times 96$ with a random rotation from $[-180^{\circ},180^{\circ}]$ in our experiments, aiming to comprehensively evaluate the performance of representative methods under different patch sizes. The parameters of rigid translation and rotation are converted to ground truth $H$ according to the mapping relationship of Equation 6. It should be noted that each pair of fingerprints comes from two impressions respectively to avoid possible information leakage caused by the same grayscale relationship or distortion pattern.
IV-B Matching Protocols
In order to balance the number of genuine / imposter matches and different fingerprint types / modalities, we implemented specific matching protocols during training and testing. For the FVC series, each fingerprint is combined with all mated fingerprints (with order) as genuine match pairs, and the first image of each pair is combined with a random fingerprint of another finger as a new imposter match pair. Taking FVC2002 DB1_A as an example, a total of $8\times 7\times 100\times 2=11,200$ image pairs are combined in this way. This protocol is also applied to NIST SD14 with a slightly different that symmetric matches are avoided, considering the large number of rolled images. All combined pairs of full fingerprints are then randomly cropped to simulate several partial fingerprint scenes, with the number of executions set to $1$ (FVC2002 DB1_A & DB3_A) and $4$ (the others). As mentioned above, some representative datasets are merged as Hybrid DB and subsequently divided, secifically $430,768$ for training and $22,672$ for testing, which are distinguished by suffixes ‘_A’ and ‘_B’ in the following experiments. The selection of genuine matches in real partial fingerprint dataset THU Small also follows this principle, while pairing the first impression of every finger with each other as imposter pairs. All detailed usage information is listed in corresponding footnotes of Table I.
V Experiments
In this section, several representative state-of-the-art fingerprint algorithms are compared with our proposed approach sufficiently and comprehensively, including:
  • A-KAZE [7], a simple but comparative partial fingerprint recognition algorithm based on universal key points;
  • VeriFinger v12.0 [10], a widely used, top performing commercial software mainly based on minutiae;
  • DeepPrint [13], a global fixed-length representation based fingerprint recognition algorithm which is highly influential;
  • LDRF [51], another region representation based fingerprint recognition algorithm that extracts more flexible localized deep descriptors;
  • PFVNet [8], a state-of-the-art partial fingerprint verification algorithm.
In particular, since the effective information of partial fingerprints is significantly reduced compared to full-size fingerprints, the absolute pose rectification pre-step adopted in DeepPrint [13] and LDRF [51] cannot be performed as expected. Therefore, in this paper we use AlignNet, part of PFVNet [8], to estimate and correct the relative pose of image pairs as substitute, which is marked by subscript ‘*’.
We conducted extensive experiments to thoroughly evaluate the performance of above algorithms under partial fingerprints at different sizes, which in terms of score distribution, matching performance, alignment accuracy, efficiency and visual analysis. In addition, exhaustive ablation studies are demonstrated to validate the effectiveness of corresponding modules and strategies proposed in this paper.
V-A Implementation Details
All partial fingerprint training processes are performed on the generated dataset Hybrid DB_A (where patch size is fixed at $160\times 160$) described in Section IV with a initial learning rate of $1\mathrm{e}{-3}$ (end of $1\mathrm{e}{-6}$), cosine annealing scheduler, default AdamW optimizer and batch size of $128$ until convergence (about $12$ epochs). The pre-training task proposed in Section III-A is trained for $300$ epochs using a subset of $5,000$ rolled images with corresponding random augmentation. On the other hand, considering that both fixed-length representation based methods [13, 51] require large amounts of training data, we completely reimplemented them on full-size fingerprints in the manner of [51]. Hybrid DB_B, which has the same source as training set, is used to test the performance under same data distribution. Besides, scenarios of other image sizes and datasets are directly used for evaluation without fine-tuning, which can also reflect the cross-domain adaptability of each model to a certain extent.
V-B Score Distribution
We first present the score distribution of genuine and imposter matches on Hybrid DB_B with image size $160\times 160$ to qualitatively demonstrate the discriminative capabilities of different algorithms. Corresponding curves are shown in Fig. 7, where the scores of A-KAZE, VeriFinger, and LDRF are linearly mapped to a numerical range that approximates $[0,1]$ for intuitive comparison. A more discrete pairwise distribution (i.e., fewer overlapping areas) indicates that the corresponding algorithm can better distinguish whether a certain fingerprint pair comes from the same finger.
Moreover, three common indicators are further computed and listed in Table II to quantitatively measure the differences between genuine and imposter probability distributions. Among them, the Jensen–Shannon divergence (JS) is defined as:
$\operatorname{KL}\left(p\|q\right)$ $=\int p(x)\log\frac{p(x)}{q(x)}dx\;,$ (11)
$\operatorname{JS}\left(p\|q\right)$ $=\frac{1}{2}\operatorname{KL}\left(p\|\frac{p+q}{2}\right)+\frac{1}{2}\operatorname{KL}\left(q\|\frac{p+q}{2}\right)\;,$
where $p$ and $q$ correspond to any two distributions. Let $\Pi(p,q)$ represent all possible joint distributions of $p$ and $q$, $F$ represent the functions of sample space, the Earth-Mover’s Distance (EMD) and Maximum Mean Discrepancy (MMD) are calculated by
$\operatorname{EMD}\left(p,q\right)=\inf_{\gamma\sim\Pi(p,q)}\mathbb{E}_{(x,y)\sim\gamma}[\|x-y\|]\;,$ (12)
and
$\operatorname{MMD}\left(F,p,q\right)=\max_{f\in F}\left(E_{x\sim p}[f(x)]-E_{y\sim q}[f(y)]\right)\;.$ (13)
Similarly, the larger these values, the farther between corresponding distribution pairs, which normally means a better identity verification performance.
Figure 7: Probability density distribution of genuine and imposter matching scores on Hybrid DB_B with image size $160\times 160$. Scales on vertical axis are hidden because we are more concerned with the relative values.
TABLE II: Probability distribution distance of genuine and imposter scores on Hybrid DB_B with image size $160\times 160$.
Method JS EMD MMD
A-KAZE [7] 0.42 0.20 0.72
VeriFinger [10] 0.70 0.59 3.21
DeepPrint* [13] 0.59 0.36 2.03
LDRF* [51] 0.23 0.08 0.20
PFVNet [8] 0.73 0.43 3.01
[…]
According to the results in Fig. 7 and Table II, our method outperforms all other methods, while VeriFinger [10] and PFVNet [8] each have sub-optimal performance in some parts. In particular, the distinguishing ability of LDRF* [51] is not ideal because it is designed based on the priori assumption of high-precision alignment (within about $9$ pixel and $4^{\circ}$ reported in original paper), which is difficult to satisfy in current partial fingerprint situations.
V-C Matching Performance
TABLE III: Matching Performance with State-of-the-Art Algorithms and Our Approach on Four Representative Datasets of Partial Fingerprint in Three Sizes.
Size Sketch Method Hybrid DB_B a FVC2002 DB1_A a FVC2002 DB3_A a THU Small b
AUC AP TAR EER AUC AP TAR EER AUC AP TAR EER AUC AP TAR EER
160 $\times$ 160 Fig. LABEL:fig:intro_method_a A-KAZE [7] 0.70 0.77 0.28 0.37 0.86 0.90 0.56 0.22 0.77 0.83 0.36 0.31 0.87 0.88 0.68 0.21
VeriFinger [10] 0.88 0.91 0.56 0.18 0.97 0.98 0.85 0.07 0.93 0.95 0.76 0.12 0.91 0.91 0.76 0.15
Fig. LABEL:fig:intro_method_b DeepPrint* [13] 0.87 0.88 0.16 0.21 0.92 0.93 0.32 0.16 0.89 0.90 0.16 0.19 0.86 0.83 0.26 0.22
LDRF* [51] 0.64 0.65 0.02 0.40 0.78 0.80 0.10 0.29 0.72 0.75 0.07 0.34 0.71 0.66 0.13 0.35
Fig. LABEL:fig:intro_method_c PFVNet [8] 0.94 0.95 0.47 0.14 0.96 0.97 0.61 0.10 0.96 0.97 0.39 0.11 0.87 0.85 0.47 0.22
[…]
128 $\times$ 128 Fig. LABEL:fig:intro_method_a A-KAZE [7] 0.63 0.68 0.15 0.42 0.73 0.80 0.32 0.34 0.67 0.72 0.14 0.39 0.74 0.74 0.40 0.34
VeriFinger [10] 0.74 0.79 0.29 0.33 0.85 0.89 0.55 0.22 0.79 0.83 0.43 0.29 0.80 0.80 0.52 0.28
Fig. LABEL:fig:intro_method_b DeepPrint* [13] 0.73 0.75 0.03 0.33 0.75 0.78 0.07 0.32 0.75 0.77 0.07 0.32 0.69 0.60 0.05 0.37
LDRF* [51] 0.59 0.60 0.01 0.44 0.63 0.65 0.02 0.41 0.63 0.65 0.03 0.41 0.60 0.50 0.03 0.43
Fig. LABEL:fig:intro_method_c PFVNet [8] 0.81 0.82 0.12 0.27 0.79 0.82 0.15 0.29 0.82 0.83 0.13 0.27 0.74 0.67 0.11 0.33
[…]
96 $\times$ 96 Fig. LABEL:fig:intro_method_a A-KAZE [7] 0.52 0.53 0.02 0.49 0.54 0.58 0.05 0.49 0.52 0.54 0.01 0.49 0.53 0.43 0.06 0.48
VeriFinger [10] 0.60 0.63 0.09 0.41 0.66 0.70 0.19 0.38 0.63 0.66 0.12 0.39 0.67 0.62 0.24 0.37
Fig. LABEL:fig:intro_method_b DeepPrint* [13] 0.64 0.65 0.01 0.40 0.61 0.64 0.02 0.42 0.64 0.65 0.01 0.40 0.56 0.44 0.01 0.45
LDRF* [51] 0.52 0.53 0.00 0.49 0.55 0.56 0.01 0.47 0.54 0.55 0.00 0.47 0.53 0.42 0.01 0.48
Fig. LABEL:fig:intro_method_c PFVNet [8] 0.71 0.71 0.03 0.36 0.65 0.67 0.04 0.40 0.69 0.70 0.03 0.38 0.64 0.52 0.02 0.42
[…]
  • a
    Image pairs are uniformly distributed within a relative translation of 0 to 100 pixels and relative rotation range of 360 degrees.
  • b
    Fingerprints are entered by volunteers in any comfortable pressing pose, with no special declaration or device restrictions.
Figure 8: Equal Error Rate matrix under different degrees of relative rigid transformation on Hybrid DB_B with image size $160\times 160$. Genuine matches are divided into 16 subsets based on the value of relative rotation $\theta$ and translation $T$, which in terms of Small, Medium, Large, and Huge segmented evenly by numerical range. Imposter matches are not modified again.
In this subsection, four classic indicators are adopted to comprehensively evaluate methods from the perspective of different deployment scenarios, including:
  • Area Under Curve (AUC) based on Receiver Operating Characteristic (ROC) curves, to evaluate the overall performance in 1:1 verification;
  • Average-Precision (AP) based on Precision-Recall (PR) curves, to approximately evaluate the search performance under 1:N closed-set;
  • True Accept Rate (TAR) under specified False Accept Rate (FAR), to rigorously evaluate the identification performance 1:N open-set;
  • Equal Error Rate (EER), a comparable standard for system trade-offs.
Table III presents the corresponding matching performance of state-of-the-art algorithms and our proposed approach on several representative datasets at different sizes. Furthermore, we group the genuine matches on Hybrid DB_B according to their relative rigid transformation relationship and compare these algorithms under different rotations and translations using EER. Fig. 8 shows the corresponding comparisons. Additionally, the Detection Error Tradeoff (DET) curves on all test databases are calculated in Fig. 9 to provide more complete information for analysis.
Figure 9: DET curves with state-of-the-art algorithms and our approach on partial fingerprints of $160\times 160$. Solid and dotted lines represent deep learning methods and traditional methods respectively.
These encouraging experimental results strongly demonstrates the superiority of our method, which outperforms previous advanced methods in almost all cases. PFVNet [8] also shows remarkable performance, especially in indicators of AUC and AP. However, it can be observed that the relative performance deteriorates markedly when training and testing data cross domains. Region representation based methods[13, 51] are inferior to others, which is reasonable because the highly generalized form of fixed-length descriptors sacrifices many potential details for faster search speed. Specifically, although more flexible localized descriptors are designed in LDRF [51], their effectiveness is limited, as analyzed in Section V-B. On the other hand, the performance of all algorithms drops significantly as the image size decreases or distance increases, but our scheme still leads and exhibits attractive robustness and stability. It is worth noting that the impact of rotation is not evident in Fig. 8, because key points based matching score do not require pose rectification and the prediction error in other algorithms remains at a similar level across different relative angles (confirmed in the following experiments).
V-D Visual Analysis
Figure 10: Representative visualization results of genuine matches based on occlusion sensitivity. Red areas indicate a large influence on matching judgment, while blue areas indicate the opposite. White numbers on the upper left are the accuracy of corresponding methods when they are verified.
Figure 11: Challenging cases where our approach fails. The way of visualization is the same as Fig. 10.
In order to provide more specific, vivid, and intuitive comparison with interpretability, we employ occlusion sensitivity [22] to approximate the contribution of distinct local areas to the overall matching judgment. Fig. 10 shows four representative visualization results. Query fingerprint in images pairs is centered and the search fingerprint is rigidly aligned based on the ground truth as the background. Subsequently, a $32\times 32$ mask is slid and applied to search fingerprints and the differences between scores before and after occlusion are calculated, which are normalized and overlaid as the heatmap of search fingerprint. Considering that there is no reference relationship between the matching scores directly obtained (or mapped) by different algorithms, we convert these scores into a comparable probability form and attached them in corresponding visualization results. Specifically, given the genuine matching score $s$ and overall distributions (shown in Fig. 7), the verification accuracy can be expressed as:
$P\left(y=1|s\right)$ (14)
$=\frac{P\left(s|y=1\right)P\left(y=1\right)}{P\left(s|y=1\right)P\left(y=1\right)+P\left(s|y=0\right)P\left(y=0\right)}$
$=\frac{f_{\mathrm{1}}\left(s\right)P\left(y=1\right)}{f_{\mathrm{1}}\left(s\right)P\left(y=1\right)+f_{\mathrm{0}}\left(s\right)P\left(y=0\right)}\;,$
where $P$ and $f$ represent the probability and probability density, $y$ represents the identity to be verified (1 is the same, 0 is different). It can be seen that algorithms based on key points and region representation expose their respective shortcomings when there is a lack of sufficient features in overlapping areas (row 1) and significant differences in ridge texture of non overlapping areas (row 2). PFVNet [8] has clear advantages in both cases, even with minimal overlap (row 3). However, it fails in the last scene with confusion in rotation. At the same time, our method demonstrates optimal stability and robustness in focusing on overlapping regions and gives correct decisions with the highest accuracy.
Meanwhile, some failure cases in Fig. 11 show that our method still needs to be improved in some extreme scenarios, such as: (i) sever defacement caused by incorrect collection conditions (row 1); (ii) significant misalignment of ridges caused by finger distortion (row 2); (iii) extremely weak texture, which refers to almost no minutiae or changes in orientation, as well as large modal differences (row 3).
V-E Alignment Accuracy
Figure 12: Estimation errors of alignment on Hybrid DB_B in image size $160\times 160$ with respective to relative rotation angle $\left|\theta\right|$ specifically, (a) rotation errors and (b) translation errors - along with relative translation difference $\left|T\right|$, which includes (c) rotation errors and (d) translation errors. Lines and corresponding light-colored bands reflect the mean and standard deviation.
The original intention of introducing the pose estimation sub stage in our network is to help it better understand the spatial contrast relationship and thus facilitate partial fingerprint verification. Nevertheless, we still assess the alignment accuracy as it is indeed a specific aspect that can be effectively compared, allowing us to gauge their capability to rectify relative positions and potential influence on subsequent matching. Similar to [8], we express the estimation error in the form of translation and rotation as:
$|\Delta\theta|$ $=|\theta^{*}-\theta|\;,$ (15)
$|\Delta T|$ $=\sqrt{(t_{\mathrm{x}}^{*}-t_{\mathrm{x}})^{2}+(t_{\mathrm{y}}^{*}-t_{\mathrm{y}})^{2}}\;,$
where symbols are defined the same as in Equation 9. As shown in Fig. 12, most methods can provide effective predictions within an acceptable range, while our approach has a slight edge. Besides, it can be seen that estimation errors of both rotation and translation increase significantly with distance and remaining roughly stable for angle. One convincing explanation is that distance greatly alters the overlap area, which drastically affects the internal feature comparison and subsequent decision-making process.
V-F Efficiency
Figure 13: Comparisons of the trade-off between performance and efficiency on Hybrid DB_B in image size $160\times 160$. The size of solid points represents the model size of deep learning methods, while hollow points represent non learning algorithms.
In this subsection, we present a comparison of the deployment efficiency across different algorithms. As can be observed from Fig 13, our proposed method achieves a new state-of-the-art trade-off between matching performance, speed, and size. This clearly highlights our streamlined joint estimation framework of pose and identity, as well as the appropriate and efficient network design. The calculated time covers a complete process from inputting an original image pair to outputting the final estimation result. All algorithms are implemented in Python and the batch size is set to 1, which are deployed on a single NVIDIA GeForce RTX 3090 GPU with a 2.4 GHz CPU.
V-G Ablation Study
Extensive ablation experiments are conducted in Table IV to examine the effects of different modules and strategies introduced in our network. Overall, the CNN based approach outperforms Vision transformer (ViT) based approach, while the CNN-Transformer hybrid series achieves the best overall performance. We attribute it to the respective advantages of convolutional and transformer blocks in extracting local features and establishing global connections. In addition, the introduction of joint estimation strategy significantly reduces the EER by 16.0%, 38.3% and 58.8% in above solutions, which proves that there is indeed an exploitable promotion relationship between pose prediction and identity verification. At the same time, the additional alignment head can be removed during deployment so that the matching speed will not be affected. Finally, the comparison between last two rows demonstrated that our proposed lightweight pre-training task can further improve the performance at a small extra training time cost.
TABLE IV: Ablation Study of the Proposed Network with Different Modules and Strategies on FVC2002 DB1_A with image size $160\times 160$.
Encoder * Decoder * + Align Head * EER
CNN ViT CNN ViT
0.25
0.21
0.47
0.29
0.17
0.07
[…]
  • §$\S$
    indicates that module has been pretrained.
    each represents specific stages in proposed network, specifically ‘Encoder’ for stage 1 & 2, ‘Decoder’ for stage 3 & 4 - C, ‘Align Head’ for stage 4 - R.
VI Conclusion
In this paper, we propose a joint estimation framework of identity verification and relative pose for partial fingerprints. A novel CNN-Transformer hybrid network is presented to combine their advantages in feature extraction and interaction, promoting the attention and utilization of valuable information. In addition, a lightweight pre-train task is designed to improve the representation ability of feature encoder by reconstructing enhanced images. Comprehensive experiments on multiple databases demonstrate the effectiveness and superiority of our proposed method. Futher studies will focus on further improving the performance on extremely low quality fingerprints, and extend the method to single-to-multiple scenarios.